# Bayesian Stability Selection and Inference on Inclusion Probabilities

### v1

@@ Abstract
Stability selection is a versatile framework for structure estimation and variable selection in high-dimensional setting, primarily grounded in frequentist principles. In this paper, we propose an enhanced methodology that integrates Bayesian analysis to refine the inference of inclusion probabilities within the stability selection framework. Traditional approaches rely on selection frequencies for decision-making, often disregarding domain-specific knowledge and failing to account for the inherent uncertainty in the variable selection process. Our methodology uses prior information to derive posterior distributions of inclusion probabilities, thereby improving both inference and decision-making. We present a two-step process for engaging with domain experts, enabling statisticians to elucidate prior distributions informed by expert knowledge while allowing experts to control the weight of their input on the final results. Using posterior distributions, we offer Bayesian credible intervals to quantify uncertainty in the variable selection process. In addition, we highlight how selection frequencies can be uninformative or even misleading when covariates are correlated with each other, and demonstrate how domain expertise can alleviate such issues. Our approach preserves the versatility of stability selection and is suitable for a broad range of structure estimation challenges.
Keywords: Bioinformatics, Bayesian Inference, Model Selection, Stability Selection, Structure Estimation, Variable Selection
@@ 1 Introduction
The estimation of discrete structures, including graphs and clusters, and the selection of variables represent long-standing and fundamental problems in statistics (Meinshausen and Bühlmann, 2010). Variable selection is a critical phase in modelling and a crucial step in data pre-processing, which involves modifying and altering raw data to make it suitable for analysis or modelling. Variable selection removes variables that are noisy, redundant, or irrelevant (Tibshirani, 1996; Müller and Welsh, 2010), which is a key challenge in modern data science (Nogueira et al., 2018).
@@
Stability selection, a widely recognised resampling-based variable selection method, determines the selection of variables based on their selection frequencies (Meinshausen and Bühlmann, 2010); which means the frequency with which variables are included in estimated models across multiple resamples.
@@
This is somewhat surprising, given that the discussion section of Meinshausen and Bühlmann (2010) includes some remarks suggesting the potential to explore stability selection from a Bayesian perspective, and despite the substantial attention stability selection has received since its inception, we have found no studies in the literature that integrate it with Bayesian analysis. In this paper, we explore how the Bayesian framework can potentially enhance stability selection by incorporating prior information, offering a comprehensive inferential tool for decision-making.
@@
We demonstrate the methodology using variable selection in the context of linear regression. However, the applications of this approach are not limited to linear regression, as stability selection itself is not limited to this problem. Although beyond the scope of this paper, our framework is applicable for more complex regression-type models.
The idea of using resamples of data to perform variable selection has been used since the past decades. For example, Chen and George (1985) leveraged bootstrap resamples (Efron, 1979) to find important prognostic factors via a stepwise regression procedure in Cox’s proportional hazards model (Cox, 1972). In this context, Rathnayake and Olive (2023) employed bootstrap sampling for variable selection in generalised linear (McCullagh, 2019) and survival models (Cox, 2018).
The stability of a variable selection method denotes the robustness of its preferences for variables across diverse training sets sampled from the same generating distribution (Kalousis et al., 2007) and has been a topic of interest in recent decades (Breiman, 1996b). Evaluation of the stability of a statistical model using random resamples has also been used for decades. For example, Altman and Andersen (1989) applied this approach to Cox’s proportional hazards regression model to identify stable variables, defined as those selected with higher frequencies across multiple resamples.
@@
In the influential work by Meinshausen and Bühlmann (2010), the innovative stability selection framework was introduced, offering a transformative approach for identifying stable variables in structure estimation tasks. The stability selection in Meinshausen and Bühlmann (2010) uses sub-samples with half-size of data to perform variable selection and selects the variables that are consistently identified as important variables across the majority of sub-samples as stable variables. Stability selection can be seen as a meta-algorithm similar to Breiman (1996a) but instead of aggregating the predictions, it aggregates the selections. Meinshausen and Bühlmann (2010) provided an asymptotic upper-bound for the per-family error rate or the expected number of falsely selected variables of the algorithm under the assumptions of exchangeability of the distribution of falsely selected variables and that the selection algorithm is not worse than random guessing.
Shah and Samworth (2013) introduced complementary pairs stability selection, a method where selection frequencies are calculated by counting the number of times a variable is selected in models fitted to complementary 50% sub-samples. Shah and Samworth (2013) introduced more rigorous upper-bounds for the expected number of variables selected with low selection probabilities.
As Holmes noted in the discussion section of Meinshausen and Bühlmann (2010), which is highly relevant to the motivation of this study, statisticians often use crisp variable subset selection to work with data owners to assess the relevance of measured variables and gain a deeper understanding of the dependence structures within the data. Holmes emphasised the importance of providing a formal mechanism for incorporating domain knowledge related to stability selection. Although the results presented by Meinshausen and Bühlmann (2010) depend solely on the data and the algorithm used, the adoption of a Bayesian framework offers a potential way to integrate background knowledge, thus enhancing the results beyond the mere data and algorithmic functions.
The principle of stability selection in Meinshausen and Bühlmann (2010) is based on the repeated sub-sampling of data, using the frequencies with which variables are selected as their selection probabilities to guide decision-making. We advocate for this approach by presenting two main arguments: (1) selection frequencies should be interpreted as sampling probabilities rather than inferential probabilities. As Clayton (2021) explained, “Sampling probabilities go from hypothesis to data: Given an assumption, what will we observe, and how often? Inferential probabilities go from data to hypothesis: Given what we observed, what can we conclude, and with what certainty?”. In this paper, we propose a procedure to enhance sampling probabilities to inferential probabilities, and (2) a sole reliance on frequencies overlooks the importance of background knowledge, thus dismissing the wealth of human understanding in favour of a purely randomised process.
In this paper, we define the inclusion probability as the probability that a variable is incorporated into the estimation of the true underlying model, taking into account both the data and the background knowledge. Our objective is to use selection frequencies to adjust prior probabilities rather than to adopt a position of complete ignorance regarding the variables involved. We further demonstrate, through simulations and analysis of real biological data sets, that correlations between covariates can lead to erratic selection frequencies. In such cases, the incorporation of prior information can improve the decision-making process.
Bayesian variable selection methods, such as Gibbs variable selection (Carlin and Chib, 1995), stochastic search variable selection (George and McCulloch, 1993), adaptive shrinkage (Xu, 2003; Figueiredo, 2003), Bayesian LASSO (Park and Casella, 2008), and reversible jump Markov chain Monte Carlo (Green, 1995), leverage prior information to elucidate parameters, including indicator variables for spike-and-slab models, as well as the location, scale, and rate parameters of coefficients, or the number of variables incorporated into the model (O’Hara and Sillanpää, 2009).
In a Bayesian setting, rarer prior probabilities necessitate greater evidence to surpass the prior and establish an alternative hypothesis. We identified a lack of Bayesian variable selection methods that effectively regulate the extent to which priors dominate the results, thus preventing deterministic outcomes. In this context, we propose a novel and straightforward method for integrating ‘domain expert’ background knowledge into the analysis process, where such expertise is available. This approach improves communication with data owners, allowing them to understand how their expertise is incorporated and the degree to which their input influences the results.
In this paper, we propose the use of the Beta distribution as the prior distribution of inclusion probabilities, which has gained increasing attention in variable selection, as demonstrated by recent studies of Liu and Ročková (2023) and Sparapani et al. (2023). In this regard, we introduce a novel method for determining the parameters of the Beta prior distribution using the experience and knowledge of experts. Specifically, we ask two key questions for each variable under selection, where the first determines the extent to which their background knowledge influences the final results, and the second assesses the perceived importance of the variables. The responses to these questions are then used to elucidate the prior distributions. Although we do not have access to domain experts for our data examples, which limits our ability to directly demonstrate the effectiveness of this approach, we use previously published results to construct informative priors. These priors serve as a proxy for expert knowledge, thereby allowing us to simulate the incorporation of expert insights into our methodology.
To the best of our knowledge, there has been no dedicated research addressing the integration of stability selection and the Bayesian framework, leaving a notable gap in the literature. This study aims to integrate stability selection, which is grounded in frequentist principles, with a novel Bayesian framework to provide inferential inclusion probabilities and Bayesian credible intervals instead of solely relying on sampling frequencies for decision-making. Furthermore, this study poses a question specifically designed to explicitly assess and regulate the degree of subjectivity involved in the final results.
@@
The structure of this paper is organised as follows. Section 2 details the proposed methodology. Section 3 describes the real and synthetic data sets used to apply the methodology. The results of the application of the method to real and synthetic data are presented in Section 4. Finally, Section 5 offers some remarks on Bayesian stability selection, while Section 6 provides a summary and concludes the paper.
@@ 2 Methodology
Stability selection, as originally formulated by Meinshausen and Bühlmann (2010), is inherently agnostic to prior information about the variables, relying solely on iterative application of a selection method to different sub-samples of observations. We propose a Bayesian extension to this approach, designed to incorporate expert knowledge into the data-driven selection process. We will outline the specifics of this extension after providing a concise overview of the foundational methodology introduced by Meinshausen and Bühlmann (2010).
@@ 2.1 Stability Selection
We consider a data set $\mathcal{D}=\{(\boldsymbol{x}^{\top}_{i},y_{i})\}_{i=1}^{n}$ comprising of a univariate response variable $y_{i}\in\mathbb{R}$ and a $p$-dimensional vector of fixed covariates $\boldsymbol{x}^{\top}_{i}\in\mathbb{R}^{p}$. Typically, when modelling pairs $(\boldsymbol{x}^{\top}_{i},y_{i})$ in the linear regression context, these are assumed to be independent and identically distributed (i.i.d.).
The linear regression model is formally defined as $\boldsymbol{Y}=\beta_{0}+X\boldsymbol{\beta}+\boldsymbol{\varepsilon},$ where $X\in\operatorname{mat}(n\times p)$ denotes the design matrix, $\boldsymbol{Y}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector representing the univariate response variable, $\beta_{0}\in\mathbb{R}$ denotes the intercept term, $\boldsymbol{\beta}\in\mathbb{R}^{p}$ denotes the $p$-dimensional vector of regression coefficients for the $p$ non-constant covariates, and $\boldsymbol{\varepsilon}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector of random errors. It is assumed that the components of $\boldsymbol{\varepsilon}$ are i.i.d. and independent of the covariates.
In variable selection, following the terminology of Meinshausen and Bühlmann (2010), it is generally assumed that the covariates can be divided into two distinct groups: the signal group $S=\{k\neq 0\mid\beta_{k}\neq 0\}$ and the noise group $N=\{k\neq 0\mid\beta_{k}=0\}$, where $S\cap N=\emptyset$. A primary goal of a variable selection method is to accurately estimate the signal group $S$.
The stability selection framework requires an appropriate selection method to be implemented within it. One widely used method for variable selection in linear regression setting is the Least Absolute Shrinkage and Selection Operator (LASSO; Tibshirani, 1996), a regularisation technique that employs the $l_{1}$-norm to perform shrinkage on vanishing coefficients. The LASSO estimator is defined as
$\hat{\beta}_{0}(\lambda),\boldsymbol{\hat{\beta}}(\lambda)=\operatorname*{arg\,min}_{\beta_{0}\in\mathbb{R},\boldsymbol{\beta}\in\mathbb{R}^{p}}\left(\|\boldsymbol{Y}-\beta_{0}-X\boldsymbol{\beta}\|_{2}^{2}+\lambda\sum_{k=1}^{p}|\beta_{k}|\right),$ (1)
where $\lambda\in\mathbb{R}^{+}$ is the LASSO regularisation parameter. The set of non-zero coefficients $\hat{S}(\lambda)=\{k\neq 0\mid\hat{\beta}_{k}(\lambda)\neq 0\}$ can be identified using the solution from Equation (1) via convex optimisation.
Meinshausen and Bühlmann (2010) defined the stable set
$\hat{S}^{\text{stable}}=\{j\mid\max_{\lambda\in\Lambda}(\hat{\Pi}_{j}^{\lambda})\geq\pi_{\text{thr}}\};\;j=1,\ldots,p,$
where $\Lambda$ denotes the set of regularisation values, $10<\pi_{\text{thr}}<$ 1 denotes a threshold for decision-making in variable selection, and $\hat{\Pi}_{j}^{\lambda}$ denotes the selection frequency of the $j$th variable given the regularisation parameter $\lambda$.
@@ 2.2 Bayesian Stability Selection
We now move forward with the introduction of Bayesian stability selection. In this study, the value $\lambda$ is selected by identifying the largest value of $\lambda$ such that the error remains within one standard error of the minimum error obtained from the cross-validation (Hastie et al., 2009) on all the data $\mathcal{D}$. We use the obtained $\lambda$ value throughout all iterations of the stability selection process. If exploration of different $\lambda$ values is desired, one may employ a grid of $\lambda$ values and repeat the process for each one. Although there are alternative approaches for variable selection in the linear regression model, exploring various methods to determine the regularisation technique and penalty term in stability selection is beyond the scope of this paper.
During each iteration of the stability selection procedure, a random sub-sample comprising half the size of the original data set $\mathcal{D}$ is used. The LASSO model is fitted to this sub-sample, and the binary selection outcomes are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the total number of sub-samples.
In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is selected as a member of the signal set $\hat{S}(\lambda)$ by applying LASSO to the $b$th sub-sample. Conversely, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is included in the noise set $\hat{N}(\lambda)$ when LASSO is applied to the $b$th sub-sample.
By assuming independence in the selection of variables, similar to Bodinier et al. (2023), we consider a Bernoulli distribution for the selection of each variable in stability selection iterations $M(\lambda)_{bj}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\text{Bernoulli}(\Pi_{j}^{\lambda});\;j=1,\ldots,p$, where $\Pi_{j}^{\lambda}$ denotes the inclusion probability of the $j$th variable given $\lambda$.
@@
$\sum_{b=1}^{B}M(\lambda)_{bj}\sim\text{Binomial}(B,\Pi_{j}^{\lambda});\quad j=1,\ldots,p.$ (2)
To go further, it is necessary to select an appropriate prior distribution for $\Pi_{j}^{\lambda}$. The Beta distribution $\text{Beta}(\alpha,\beta)$ is particularly suitable for modelling probabilities and has support $[0,1]$. Moreover, when the shape parameters of the Beta distribution are not less than $1$, that is, $\alpha,\beta\geq 1$, they are easily interpretable, facilitating the incorporation of background knowledge by directly translating it into a statistical object.
In the Beta distribution, the shape parameters $\alpha$ and $\beta$ offer significant insight into the form and characteristics of the distribution. When $\alpha,\beta>1$, the relative magnitudes of $\alpha$ and $\beta$ determine the skewness of the distribution. Specifically, if $\alpha>\beta$, the corresponding density is left skewed, with the mode shifted toward $1$. In contrast, if $\beta>\alpha$, the density is right skewed, with the mode shifted toward 00. When $\alpha=\beta$, the Beta distribution is symmetric. Importantly, the Beta distribution encompasses the standard Uniform distribution $\mathcal{U}(0,1)$ as a special case when $\alpha=\beta=1$, which represents a non-informative flat prior over the interval $[0,1]$.
Therefore, expert knowledge can be incorporated into the Beta distribution by expressing beliefs through an according choice of the shape parameters $\alpha$ and $\beta$, thereby incorporating prior information into the distribution in a straightforward manner. We will further elaborate on this below.
An additional advantage of the Beta distribution is its conjugacy with the Binomial distribution, which allows for the straightforward estimation of posterior distributions through reparametrisation, thereby obviating the need for complex numerical methods in posterior estimation.
By employing the Beta distribution as prior distribution of inclusion probability for the $j$th variable we have
$\pi(\Pi_{j}^{\lambda})=f(\Pi_{j}^{\lambda};\alpha_{j},\beta_{j})=\frac{({\Pi_{j}^{\lambda}})^{\alpha_{j}-1}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1}}{B(\alpha_{j},\beta_{j})};\quad j=1,\ldots,p\quad\text{and}\quad\alpha_{j},\beta_{j}\geq 1,$ (3)
@@
Given $n_{j}$ selections of the $j$th variable out of the $B$ iterations, the likelihood function for $\Pi_{j}^{\lambda}$ based on Equation (2) is
$\mathcal{L}(\Pi_{j}^{\lambda}|M(\lambda))=\binom{B}{n_{j}}(\Pi_{j}^{\lambda})^{n_{j}}(1-\Pi_{j}^{\lambda})^{B-n_{j}};\quad j=1,\ldots,p.$ (4)
@@
The posterior distributions can be obtained using the prior distributions in the form of Equation (3) and the information of the rows as observed samples using Equation (4). In other words, each row of $M(\lambda)$ is considered an independent experiment and the prior distribution will be updated in light of the information of the rows. Using Bayes’ theorem, the posterior distribution is proportional to the product of the prior distribution and the likelihood function (Jeffreys, 1934), that is
$\pi(\Pi_{j}^{\lambda}|M(\lambda))\propto(\Pi_{j}^{\lambda})^{\alpha_{j}-1+n_{j}}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1+B-n_{j}},$
@@
which is of the form of the Beta distribution $\Pi_{j}^{\lambda}|M(\lambda)\sim\text{Beta}(\alpha^{\prime}_{j}=\alpha_{j}+n_{j},\beta^{\prime}_{j}=\beta_{j}+B-n_{j})$. Therefore, the posterior density function is as follows
$\pi(\Pi_{j}^{\lambda}|M(\lambda))=f(\Pi_{j}^{\lambda};\alpha_{j},\beta_{j},n_{j},B)=\frac{(\Pi_{j}^{\lambda})^{\alpha_{j}+n_{j}-1}(1-\Pi_{j}^{\lambda})^{\beta_{j}+B-n_{j}-1}}{B(\alpha_{j}+n_{j},\beta_{j}+B-n_{j})}.$ (5)
Equation (5) provides the posterior probability distribution for the inclusion probabilities of variables, allowing for the evaluation of the relevance of variables based on the characteristics of the distribution. The posterior distributions can be used to construct non-asymptotic Bayesian credible intervals for inclusion probabilities, into which the true inclusion probabilities fall into them with the desired probability; providing a clear interpretable tool for communicating the results. In addition, density plots and other distributional characteristics, such as variance, skewness, and kurtosis, can offer a clearer understanding of the selection status of variables.
2.3 Selecting non-informative and informative priors

### v2

# Bayesian Stability Selection and Inference on Selection Probabilities

@@ Abstract
Stability selection is a versatile framework for structure estimation and variable selection in high-dimensional setting, primarily grounded in frequentist principles. In this paper, we propose an enhanced methodology that integrates Bayesian analysis to refine the inference of selection probabilities within the stability selection framework. Traditional approaches rely on selection frequencies for decision-making, often disregarding domain-specific knowledge. Our methodology uses prior information to derive posterior distributions of selection probabilities, thereby improving both inference and decision-making. We present a two-step process for engaging with domain experts, enabling statisticians to construct prior distributions informed by expert knowledge while allowing experts to control the weight of their input on the final results. Using posterior distributions, we offer Bayesian credible intervals to quantify uncertainty in the variable selection process. Furthermore, we demonstrate how the integration of prior knowledge reduces the variance of selection probabilities, thereby improving the stability of decision-making. Our approach preserves the versatility of stability selection and is suitable for a broad range of structure estimation challenges.
Keywords: Bioinformatics, Bayesian Inference, Feature Selection, Prior Elicitation, Structure Estimation, Variable Selection
@@ 1 Introduction
The estimation of discrete structures, including graphs and clusters, and the selection of variables represent long-standing and fundamental problems in statistics (Meinshausen and Bühlmann, 2010). Stability selection, a widely recognized resampling-based variable selection framework, determines the selection of variables based on their selection frequencies (Meinshausen and Bühlmann, 2010); which means the frequency with which variables are included in estimated models across multiple random resamples of observations. Stability selection has remained an active area of research since its inception (see, for example, Bodinier et al. (2023) and Kissel and Mentch (2024)).
@@
This is somewhat surprising, given that the discussion section of Meinshausen and Bühlmann (2010) includes some remarks suggesting the potential to explore stability selection from a Bayesian perspective, and despite the substantial attention stability selection has received since its inception, we have found no studies in the literature that integrate it with Bayesian analysis. In this paper, we explore how the Bayesian framework can enhance stability selection by incorporating prior information, offering a comprehensive tool for decision-making and inference.
@@
We demonstrate the methodology using variable selection in the context of linear regression. However, the applications of this approach are not limited to linear regression, as stability selection itself is not limited to this problem.
@@
In the influential work by Meinshausen and Bühlmann (2010), the stability selection framework was introduced, offering a transformative approach for identifying stable variables in structure estimation tasks. The stability selection in Meinshausen and Bühlmann (2010) uses sub-samples of observations with half-size of data to perform variable selection and selects the variables that are consistently identified as important variables across the majority of sub-samples as stable variables. Later, Shah and Samworth (2013) introduced complementary pairs stability selection, a method where selection frequencies are calculated by counting the number of times a variable is selected in models fitted to complementary 50% sub-samples.
Meinshausen and Bühlmann (2010) and Shah and Samworth (2013) defined selection probabilities solely based on selection frequencies, interpreting these as observed proportions. Although stability selection has shown significant utility across a diverse range of applications, as highlighted by Lindley (1975) and Jaynes (2003), the foundation of probability lies within the domain of human knowledge, while frequency is rooted in the external world. We define the selection probability as the probability that a variable is relevant to the response variable through a fixed model, taking into account both data and background knowledge. Our objective is to use selection frequencies to adjust prior probabilities. In addition, we examine a non-informative prior for selection probabilities, using it as a proxy for frequentist selection probabilities. We then compare the variance of posterior selection probabilities with and without prior knowledge, demonstrating how the incorporation of prior knowledge reduces the variance, thereby enhancing the stability of decision-making.
In this context, we introduce a novel yet intuitive approach for integrating domain expert knowledge into the selection process when such expertise is available, while explicitly managing the inherent subjectivity of the final outcomes. This method facilitates effective communication with data owners by providing transparency regarding how their expertise is incorporated into the analysis and the extent of its influence on the results.
@@
The structure of this paper is organized as follows. Section 2 details the proposed methodology. Section 3 describes the real and synthetic datasets used to apply the methodology. The results of the application of the method to real and synthetic data are presented in Section 4. Section 5 offers some remarks on Bayesian stability selection, and finally, Section 6 provides a summary and concludes the paper.
@@ 2 Methodology
We consider a dataset $\mathcal{D}=\{(\boldsymbol{x}^{\top}_{i},y_{i})\}_{i=1}^{n}$ comprising of a univariate response variable $y_{i}\in\mathbb{R}$ and a $p$-dimensional vector of fixed covariates $\boldsymbol{x}^{\top}_{i}\in\mathbb{R}^{p}$. We assume that the pairs $(\boldsymbol{x}^{\top}_{i},y_{i})$ are independent and identically distributed (i.i.d.).
The linear regression model is formally defined as $\boldsymbol{Y}=\boldsymbol{\beta_{0}}+X\boldsymbol{\beta}+\boldsymbol{\varepsilon},$ where $X\in\operatorname{mat}(n\times p)$ denotes the design matrix, $\boldsymbol{Y}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector representing the univariate response variable, $\boldsymbol{\beta_{0}}\in\mathbb{R}^{n}$ denotes the constant $n$-dimensional vector of intercept term, $\boldsymbol{\beta}\in\mathbb{R}^{p}$ denotes the $p$-dimensional vector of regression coefficients for the $p$ non-constant covariates, and $\boldsymbol{\varepsilon}\in\mathbb{R}^{n}$ denotes the $n$-dimensional vector of random errors.
It is assumed that $E(\boldsymbol{Y}\mid X)$ is linear in parameters, $E(\boldsymbol{\varepsilon}\mid X)=0$, $\text{Var}(\boldsymbol{\varepsilon}\mid X)=\sigma^{2}<\infty$, and $\text{Cov}(\boldsymbol{\varepsilon},X_{j})=0$, for $j\in\{1,2,\dots,p\}$, where $X_{j}$ denotes the $j$th column of $X$. It is also assumed that the columns of $X$ are not linearly dependent of each other, and the components of $\boldsymbol{\varepsilon}$ are independent of each other.
@@ 2.1 Stability Selection
In variable selection, following the terminology of Meinshausen and Bühlmann (2010), it is generally assumed that the covariates can be divided into two distinct groups: the signal group $S:=\{k\neq 0\mid\beta_{k}\neq 0\}$ and the noise group $N:=\{k\neq 0\mid\beta_{k}=0\}$, where $S\cap N=\emptyset$. A primary goal of a variable selection algorithms is to accurately estimate the signal group $S$.
The stability selection framework requires an appropriate selection algorithm to be implemented within it. One widely used method for variable selection in linear regression setting is the Least Absolute Shrinkage and Selection Operator (Lasso; Tibshirani, 1996), a regularization technique that employs the $\ell_{1}$-norm to perform shrinkage on vanishing coefficients. The Lasso estimator is defined as
$\boldsymbol{\hat{\beta}_{0}}(\lambda),\boldsymbol{\hat{\beta}}(\lambda)=\operatorname*{arg\,min}_{\boldsymbol{\beta_{0}}\in\mathbb{R}^{n},\boldsymbol{\beta}\in\mathbb{R}^{p}}\left(\|\boldsymbol{Y}-\boldsymbol{\beta_{0}}-X\boldsymbol{\beta}\|_{2}^{2}+\lambda\sum_{k=1}^{p}|\beta_{k}|\right),$ (1)
where $\lambda\in\mathbb{R}^{+}$ is the Lasso regularization parameter. The set of non-zero coefficients $\hat{S}(\lambda):=\{k\neq 0\mid\hat{\beta}_{k}(\lambda)\neq 0\}$ can be identified using the solution from Equation (1) via convex optimization.
Meinshausen and Bühlmann (2010) proposed applying the selection algorithm (here, Lasso) on the different random sub-samples of observations and defined the stable set $\hat{S}^{\text{stable}}:=\{j\mid\max_{\lambda\in\Lambda}(f_{j}^{\lambda})\geq\pi_{\text{thr}}\};\;j=1,\ldots,p$, where $\Lambda$ denotes the set of regularization values, $10<\pi_{\text{thr}}<$ 1 denotes a threshold for decision-making in variable selection, and $f_{j}^{\lambda}$ denotes the selection frequency of the $j$th variable given the regularization parameter $\lambda$.
@@ 2.2 Bayesian Stability Selection
In this paper, we conceptualize stability selection as a repetitive experiment. Given that one of the primary objectives of this paper is to perform inference on selection probabilities, it is essential to carefully examine both the systematic and random components of this experiment. To ensure that the variability in the results of stability selection across different iterations arises solely from changes in the training dataset, and without loss of generality, we restrict the hyper-parameter set $\Lambda$ to a single value, denoted as $\lambda$, which is consistently applied across all iterations of the stability selection procedure.
The value $\lambda$ is selected by identifying the largest value of $\lambda$ such that the prediction error remains within one standard error of the minimum error obtained from the cross-validation (Hastie et al., 2009) on all the data $\mathcal{D}$. If exploration of different $\lambda$ values is desired, one may employ a grid of $\lambda$ values and repeat the process for each one.
During each iteration of the stability selection procedure, a random sub-sample comprising half the size of the original dataset $\mathcal{D}$ is used. The Lasso model is fitted to the current sub-sample, and the binary selection results are recorded as a row in the binary selection matrix $M(\lambda)\in\operatorname{mat}(B\times p)$, where $B$ denotes the total number of sub-samples. In this context, $M(\lambda)_{bj}=1$ indicates that the $j$th variable is selected as a member of the signal set $\hat{S}(\lambda)$ by applying Lasso to the $b$th sub-sample. In contrast, $M(\lambda)_{bj}=0$ signifies that the $j$th variable is included in the noise set $\hat{N}(\lambda)$ when Lasso is applied to the $b$th sub-sample.
By assuming independence in the selection of variables, similar to Bodinier et al. (2023), we consider a Bernoulli distribution for the selection of each variable in stability selection iterations $M(\lambda)_{bj}\stackrel{{\scriptstyle\text{ind}}}{{\sim}}\text{Bernoulli}(\Pi_{j}^{\lambda});\;j=1,\ldots,p$, where $\Pi_{j}^{\lambda}$ denotes the selection probability of the $j$th variable given $\lambda$ that is, $P(j\in S(\lambda)\mid\lambda,K)$ where $K$ denotes the expert’s prior knowledge about the relevance of the covariate.
@@
$n_{j}:=\sum_{b=1}^{B}M(\lambda)_{bj}\sim\text{Binomial}(B,\Pi_{j}^{\lambda});\quad j=1,\ldots,p.$ (2)
To go further, it is necessary to select an appropriate prior distribution for $\Pi_{j}^{\lambda}$. The Beta distribution $\text{Beta}(\alpha,\beta)$ has been of great interest in modeling selection probabilities (e.g., see Kohn et al. (2001); Bottolo and Richardson (2010); Liang et al. (2018); Staerk et al. (2024)) and has support $[0,1]$. Moreover, when the shape parameters of the Beta distribution are greater than or equal to $1$, that is, $\alpha,\beta\geq 1$, they are easily interpretable, facilitating the incorporation of background knowledge by directly translating it into a statistical object.
In literature, several approaches have been proposed to incorporate prior knowledge into the prior distributions of selection probabilities when using the Beta distribution. Kohn et al. (2001) suggested fixing the expectation and variance of the prior distribution based on available background knowledge and determining the parameters of the distribution by solving two simultaneous linear equations. Hu and Johnson (2009) proposed the prior distribution $\text{Beta}(\delta p,(1-\delta)p)$, where $10<\delta<$ 1 is termed the precision parameter. This approach is designed to maintain the influence of the prior in high-dimensional settings and discourage large model sizes. Ley and Steel (2009) proposed fixing $\alpha=1$ and determining $\beta$ by solving the corresponding equations after specifying the expectation and variance of the distribution based on the background knowledge. Castillo and Van Der Vaart (2012) demonstrated that the use of small $\alpha$ and large $\beta$ is more effective in high-dimensional settings, where the number of covariates significantly exceeds the number of observations. They recommended setting $\alpha=1$ and $\beta=p$, to promote sparsity in the model. Lastly, Liang et al. (2018) recommended the use of empirical Bayes methods, as described in Carlin and Louis (1996), to determine the parameters of the Beta distribution.
In the Beta distribution, the shape parameters $\alpha$ and $\beta$ offer significant insight into the form and characteristics of the distribution. When $\alpha,\beta>1$, the relative magnitudes of $\alpha$ and $\beta$ determine the skewness of the distribution. Specifically, if $\alpha>\beta$, the corresponding density is left skewed, with the mode shifted toward $1$. In contrast, if $\beta>\alpha$, the density is right skewed, with the mode shifted toward 00. When $\alpha=\beta$, the Beta distribution is symmetric. Importantly, the Beta distribution encompasses the standard Uniform distribution $\mathcal{U}(0,1)$ as a special case when $\alpha=\beta=1$, which represents a non-informative flat prior over the interval $[0,1]$. Therefore, expert knowledge can be incorporated into the Beta distribution by expressing beliefs through an according choice of the shape parameters $\alpha$ and $\beta$, thereby incorporating prior information into the distribution in a straightforward manner. We will further elaborate on this in Section 2.3.
An additional advantage of the Beta distribution is its conjugacy with the Binomial distribution, which allows for the straightforward estimation of posterior distributions through reparametrization, thereby obviating the need for complex numerical methods in posterior estimation.
By employing the Beta distribution as prior distribution of selection probability for the $j$th variable we have
$\pi(\Pi_{j}^{\lambda}\mid K)=\frac{({\Pi_{j}^{\lambda}})^{\alpha_{j}-1}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1}}{B(\alpha_{j},\beta_{j})};\quad j=1,\ldots,p\quad\text{and}\quad\alpha_{j},\beta_{j}\geq 1,$ (3)
@@
Given $\Pi_{j}^{\lambda}$, the probability of observing the data which is referred to as the likelihood function (Jeffreys, 1934) based on Equation (2) is
$\mathcal{L}(n_{j}\mid\Pi_{j}^{\lambda},K)=\binom{B}{n_{j}}(\Pi_{j}^{\lambda})^{n_{j}}(1-\Pi_{j}^{\lambda})^{B-n_{j}};\quad j=1,\ldots,p.$ (4)
@@
The posterior distributions can be obtained using the prior distributions in the form of Equation (3) and the information of the rows as observed samples using Equation (4). In other words, each row of $M(\lambda)$ is considered an independent experiment and the prior distribution will be updated in light of the information of the rows. Using Bayes’ theorem (Bayes and Price, 1763), the posterior distribution is proportional to the product of the prior distribution and the likelihood function (Jeffreys, 1934), that is
$\pi(\Pi_{j}^{\lambda}\mid n_{j},K)\propto(\Pi_{j}^{\lambda})^{\alpha_{j}-1+n_{j}}(1-\Pi_{j}^{\lambda})^{\beta_{j}-1+B-n_{j}},$
@@
which is of the form of the Beta distribution $\Pi_{j}^{\lambda}\mid n_{j},K\sim\text{Beta}(\alpha^{\prime}_{j}=\alpha_{j}+n_{j},\beta^{\prime}_{j}=\beta_{j}+B-n_{j})$.
The posterior distribution allows for the evaluation of the relevance of variables based on the characteristics of the distribution. In this paper, the expectation values of the posterior distributions of selection probabilities are used as a benchmark for decision-making; however, other approaches such as the one proposed by Barbieri and Berger (2004) are also applicable.
2.3 Assigning non-informative and informative priors